# Factual Knowledge in Language Models: Robustness and Anomalies under Simple Temporal Context Variations

### v3

@@ Abstract
This paper explores the robustness of language models (LMs) to variations in the temporal context within factual knowledge. It examines whether LMs can correctly associate a temporal context with a past fact valid over a defined period, by asking them to differentiate correct from incorrect contexts. The accuracy of LMs is analyzed along two dimensions: the distance of the incorrect context from the validity period and the granularity of the context. To this end, a dataset called TimeStress is introduced, enabling the evaluation of 18 diverse LMs. Results reveal that the best LM achieves perfect accuracy for only 6% of the studied facts, with critical errors that humans would not make. This work highlights the limitations of current LMs in temporal representation. We provide all data and code for further research [^1] Github URL will be added after publication.
@@ 1 Introduction
When a Language Model (LM) completes the textual prompt "The capital of France is" with "Paris", it demonstrates that it has stored this fact somewhere in its parameters. However, as shown by numerous studies Elazar et al. (2021); Dong et al. (2023); Hagen et al. (2024); Kassner and Schütze (2020), this type of factual knowledge is not necessarily robust to certain variations in the prompt (use of paraphrases, aliases, typographical errors, negations, etc.). Among these variability factors, the temporal dimension of factual knowledge has been less studied. Thus, in this paper, we study the robustness of LMs’ factual knowledge in the face of variations in the temporal context.
While the state of the art has demonstrated certain biases in LMs related to the temporal distribution of their training data or their weakness in reasoning with temporal concepts, our work aims to quantify how well LMs can correctly associate a temporal context (e.g., a year, a date) with a past fact, that is, a fact with a certain period of validity. More specifically, the research questions addressed are:
@@
    Do LMs distinguish correct and incorrect temporal contexts for facts?
@@
The First Workshop on Large Language Model Memorization (L2M2)
Figure 1: The robustness of the LM on a fact is evaluated by asking it to differentiate a set of correct and incorrect statements. The temporal context is varied along two dimensions: its position on the timeline (columns 1 and 2) and its granularity (columns 1 and 3).
@@
  • The release of a dataset, TimeStress, consisting of popular factual knowledge, temporally annotated and verbalized. This dataset allows for the replication of our experiments but also opens avenues for other studies on temporality.
  • Highlighting the low robustness of current LMs regarding their factual knowledge when it comes to positioning them in time, as well as critical errors that a human would not make. These results reveal the shortcomings of LMs in terms of internal representation of temporality, including for large models (18 models tested across various sizes and families).
In the following sections, we first discuss related work (Section 2). Then, we elaborate on the paper’s issues and present the TimeStress dataset (Section 3). Finally, we describe our experiments and analyze their results (Section 4). The source code and data to reproduce our results will be published soon.
@@ 2 Related work
Several studies have examined the temporal reasoning capabilities of LMs Zhang and Choi (2021); Chu et al. (2024); Wei et al. (2023); Fatemi et al. (2025); Dhingra et al. (2022); Xiong et al. (2024); Su et al. (2024). Notably, the works of Chen et al. (2021) and Tan et al. (2023) each proposed a dataset in which LMs are invited to answer questions involving the understanding of the temporality of facts. While these studies share similarities with ours in terms of data (temporally annotated facts), their objectives and methodologies differ. These studies test the mastery of certain temporal logic operators (date calculations, comparisons, etc.) and evaluate an average performance of LMs based on a one-test-per-fact principle. In contrast, we focus not on reasoning ability but on robustness of knowledge, that is, the ability of an LM to never make an error when the same fact is presented with different temporal contexts.
@@ 3.1 Problem Statement
Considering a temporal fact $f=(s,r,o,a,b)$, a temporal context $\tau$ is said to be correct for $f$ if $\tau$ is fully included in $[a,b]$ (i.e., $\tau\subseteq[a,b]$), incorrect if it is not included at all ($\tau\cap[a,b]=\emptyset$), or transitional otherwise ($\tau\cap[a,b]\neq\emptyset$ and $\tau\not\subseteq[a,b]$).
@@
We introduce two metrics. Given a fact $f$ and a model $M$, we express the results using a win rate $\mathcal{W}(M,f)\in[0,1]$ of $M$ for $f$, which is the ratio of the number of times the model preferred a correct context over an incorrect context for the single fact $f$ to the number of tests performed. Additionally, a robustness metric, denoted $\mathcal{R}(M,f)$, verifies that correct contexts consistently outperform incorrect ones, defined as: $\mathcal{R}(M,f)=\mathbbm{1}[\mathcal{W}(M,f)=1]$.
For a set of facts $\{f_{1},\dots,f_{N}\}$, the metrics reported in the tables are thus the average win rates $\mathcal{W}(M)=\frac{1}{N}\sum_{f_{i}}\mathcal{W}(M,f_{i})$ and the average robustness $\mathcal{R}(M)=\frac{1}{N}\sum_{f_{i}}\mathcal{R}(M,f_{i})$. For segmentation purposes in the analyses, these global metrics can be restricted to tests conducted with temporal contexts of a specific granularity (Y, YM, or YMD).
@@ 3.2 The TimeStress Dataset
We present the TimeStress dataset, which enables our study. This dataset contains over 521,000 statements generated from 2,003 temporal facts, covering 1,883 unique entities (1,385 unique subjects and 1,113 unique objects) and 86 relations. A brief sample is provided in Table 3.2.3.
On average, each fact is associated with $11$ correct temporal contexts and $74$ incorrect ones, distributed across the three granularities Y, YM, and YMD. Each fact has enough correct and incorrect contexts to be nearly impossible for a random model to know this fact robustly by chance.
@@ 3.2.1 Quintuplet Collection
The quintuplet collection process begins with a preprocessed version of Wikidata provided in Ammar Khodja et al. (2025). This source also provides a measure of each entity’s popularity, defined as the median number of human visits per month to the Wikipedia article associated with the entity in 2020. This measure is used to define the popularity of a quintuplet, calculated as the geometric mean of the popularity of its object and subject. The relevance of this measure is supported by the observation that the likelihood of LMs knowing a fact increases with the popularity of that fact (see Appendix A.1).
We collect and filter Wikidata facts following this procedure:
  1. 1.
    All quintuplets with a validity period (i.e., a start or end date are mentioned) and whose objects are not literals, such as quantities and dates, are collected.
  2. 2.
    Quintuplets valid within two distinct periods are removed to simplify result analysis, as this allows all dates outside the validity period to be considered incorrect.
  3. 3.
    Quintuplets without a delimited validity period (i.e., a start AND end date are mentioned) are removed.
  4. 4.
    Only quintuplets that were valid prior to 2021 are retained, as it ensures that all these quintuplets are past facts for all studied LMs
  5. 5.
    Only the quintuplets that are valid for longer than three years are retained, to ensure a minimal number of correct temporal contexts of Y granularity.
We then keep only the most popular quintuplets. This results in a set of 2,098 popular quintuplets with a varied set of 86 relations.
@@ 3.2.2 Quintuplet Verbalization
The process of generating statements from quintuplets is carried out using GPT-4o. First, a prompt instructs GPT-4o to generate four linguistically diverse questions from a given tuple (subject, relation, object, year), with the following guidelines: the question must be in the past tense, begin with “In [YEAR],”, be stated in a simple and concise manner without any detail that could give clues on the answer. It should be directly followed by the answer, which is the object.
A random sample of 50 questions was manually evaluated to ensure the overall quality of the generated questions. The evaluation revealed that only 1 out of 50 questions was incorrect, while the remaining questions were perfectly constructed (Wilson confidence interval at 95% = [0.85, 0.99]), which demonstrates the high quality of the questions in our dataset.
@@ 3.2.3 Context Sampling
Based on a validity interval $[a,b]$ expressed in years, centered on $m=\frac{a+b}{2}$ and of duration $d=b-a$, temporal contexts at the year granularity are uniformly sampled over the wider interval $[m-5d,m+5d]$ with a step of $0.05\times d$. From these Y-granularity contexts, YM-granularity contexts are generated by randomly selecting a month. Similarly, YMD-granularity contexts are determined by choosing a random day from each YM-granularity context. This process creates a hierarchy among contexts derived from the same year for a given fact. Note that when a date $d_{2}$ is chosen from a higher-granularity date $d_{1}$, it is necessarily correct (or incorrect) if $d_{1}$ is. However, $d_{1}$ may be transitional while $d_{2}$ is correct or incorrect. In such cases, $d_{2}$ is excluded from the set of correct or incorrect dates. This ensures that the number of correct and incorrect dates does not vary by granularity, avoiding bias when comparing model robustness across granularities.
The facts are verbalized in natural language as question-answer pairs using GPT-4o (one question per fact). These statements are manually verified on a random sample, confirming their high quality, including the absence of syntactic or typographical errors and consistent use of the past tense. To produce the statements used in the matches, each question is prefixed with a temporal context associated with the fact.
@@ 4 Experimentation
This section details our experiments on the TimeStress dataset. As a reminder, our objectives are, in order, to measure the ability of models to distinguish correct and incorrect temporal contexts, analyze their robustness and searching for anomalies, for this task when incorrect contexts are closer or farther from the validity interval, and as the granularity of contexts becomes finer.
Numerous models from different families and sizes were tested:
  • Mistral: Mistral-Nemo-Base-2407, Mistral-7B-v0.3 Jiang et al. (2023);
  • OpenEML: OpenEML-{450M, 3B} Mehta et al. (2024);
  • Gemma2: gemma-2-{2b, 9b, 27b} Team et al. (2024);
  • Llama3.1: Llama-3.1-{8B, 70B} Grattafiori et al. (2024).
For each, both pretrained and instruction-tuned versions were considered, for a total of 18 LMs. All models were sourced from huggingface.co.
In the first series of experiments, the statements were passed to the models as plain text rather than instructions to ensure a fair comparison between pretrained and instruction-tuned models. The use of an "instruction/message" format is explored in a second phase.
@@ 4.1 Overall Mastery of Temporal Contexts
Figure 2(a) shows the average win rate for the facts in TimeStress for the top 5 models and for each temporal granularity Y, YM, and YMD, as well as for their union.
Overall, the results show that these top 5 LMs generally distinguish correct statements from incorrect ones well, with win rates ranging from 72% to 83%. Among our other findings, we observed that even smaller models (<500M parameters) perform better than chance, and the win rate logically improves with model size (Appendix E), with the best model being the largest, Llama-3.1-70B-Instruct.
Figure 3 provides a more detailed analysis by reporting the average $\log\Pr(o|f,\tau)$ as a function of the value $\alpha$, which quantifies the relative distance of $\tau$ from the validity period of $f$ (see Section 3.1).
The average is calculated across all facts, for contexts at the year granularity, and across all 18 studied LMs. We observe that the highest probabilities correspond to contexts within the validity interval ($\alpha\in[-0.5,0.5]$), while outside this interval, probabilities gradually decrease as $|\alpha|$ increases.
Finally, we note that the probability assigned to transitional contexts (here, years that are neither fully correct nor fully incorrect) is significantly higher than for incorrect contexts.
We explain this phenomenon with the following hypothesis: in the training data of LMs, transitional years are more often associated with the considered fact than other years within the validity period, as they correspond to key events such as the beginning and end of the fact (e.g., the start or end year of a presidential term).
@@
Figure 2: Average metrics on the TimeStress dataset for all studied LMs with 95% confidence intervals. The 5 most robust models are starred (*).
@@ 4.2 Robustness and Anomalies
Figure 2(b) shows the average robustness of the top 5 models across all facts in TimeStress. As a reminder, this metric is stricter and does not tolerate any error during matches for a given fact.
The scores are generally low, indicating that win rates per fact rarely reach 100%. Interestingly, the most robust model is not the one with the highest win rate, likely reflecting the importance of the training protocol in improving temporal handling. The most robust model, gemma-2-27b-it, achieves an $\mathcal{R}$ value of only about 10% for the coarsest granularity Y. This score drops to 6% when all granularities are considered. Most other models do not exceed a global robustness score of 2%.
Among our other findings, we also observed that instruction-tuned models generally outperform their pretrained counterparts. The most significant gap between the two is observed with the Llama-3.1-70B-Instruct model, which is 9.2×9.2$\times$9.2 × more robust than Llama-3.1-70B. Again, this suggests that training data and possibly the training procedure play a significant role in temporal robustness.
Finally, early signs of failure in knowledge transfer between granularities are evident due to the substantial gap between individual robustness scores for granularities and the global score. This issue is explored in more detail in this section.
Figure 3: Evolution of $\log\Pr(o|f,\tau)$ with respect to the relative distance $\alpha$, averaged across all facts in TimeStress and all LMs, for granularity Y. The number of points used to compute each bar is indicated above it.
$|\alpha|$ Plain Text Instruction
$\geq 1$ 0.28 (±0.01) 0.34 (±0.02)
$\geq 2$ 0.13 (±0.01) 0.19 (±0.02)
$\geq 3$ 0.09 (±0.01) 0.12 (±0.02)
$\geq 4$ 0.05 (±0.01) 0.07 (±0.01)
Figure 4: Proportion of incorrect dates favored over correct dates beyond a relative distance $\alpha$, when the win rate exceeds 95%.
@@
Table 4 further investigates the impact of the relative positions of incorrect contexts for granularity Y, focusing on cases where incorrect contexts cause an LM to fail in a match for facts that seem well-known to the LM, as indicated by a very high win rate ($\mathcal{W}\geq$ 95%). For now, only the "Plain Text" column is of interest. The table reveals that these incorrect contexts are not entirely concentrated around the validity period, as might reasonably be expected. Instead, a significant proportion is located far from it.
Specifically, LMs fail to achieve robustness due to dates with a distance of $|\alpha|\geq 1$ in 28% of cases. This proportion decreases to 9% for $|\alpha|\geq 3$, which remains significant given the proximity of the win rate to 100% for the facts observed here. We conducted the same analysis using win rate thresholds higher than 95% (see Appendix B). As the threshold approaches 100%, vulnerability to "easy" incorrect dates gradually decreases but never completely disappears, even when the win rate threshold is 99% and $|\alpha|\geq 4$. We conclude that this vulnerability is inherent to current LMs. While the probabilistic nature of these models may provide a tangible explanation, this behavior is clearly undesirable, as these are typically errors that a human would not make when aware of a fact’s validity period.
@@
So far in our experiments, all models have been fed statements in plain text rather than instructions. Since the performance of instruction-tuned LMs might have been underestimated, win rates and robustness scores were recalculated using an "instruction/message" format [^2] This involves constructing messages and injecting them into the chat template of each LM. Messages are constructed as in the following example: {user: "In 2011, who was the president of the USA?", assistant: "Barack Obama"}..
Figure 5 compares robustness scores calculated for the two formats. On average, robustness decreases with the use of the "instruction" format (notably for gemma-2 models), and global robustness scores remain very low. However, no clear conclusions emerge regarding the positive or negative impact of this format, as the effect varies significantly across models.
Next, the "Instruction" column of Table 4 complements our previous analysis on the impact of the relative position of incorrect contexts for high win-rate facts. This time, the "instruction" format degrades performance with more critical errors (i.e., far from the validity period). These differences are statistically significant, except for $\alpha\geq 4$.
@@
Figure 5: Average $\mathcal{R}$ across all granularities for facts in TimeStress based on the format of statements submitted to the models: plain text (blue) or instruction (orange).
@@
Figure 6: Average success rate of knowledge transfer between granularity pairs for the 5 most robust LMs with queries in plain text (top) or instructions (bottom). Wilson confidence intervals at 95% are shown.
LMs fail to propagate their knowledge across granularities.
@@
We examine the ability of LMs to propagate knowledge of a fact across different temporal granularities. TimeStress allows comparisons between two granularities because the three studied granularities have the same number of correct and incorrect dates for all temporal facts. The only difference between two granularities is the addition of a random month and/or day, which does not affect validity when transitioning from a lower granularity to a higher granularity (e.g., from Y to YM). If a fact is incorrect for an entire year, it remains incorrect for any month or date within that year.
@@
We consider a fact $f$ to be known for a granularity by a model $M$ if $\mathcal{R}(M,f)=1$. This definition can apply to a given granularity. For example, a fact is known at the Y granularity if all matches with temporal contexts at the year granularity were won. For each of the 5 most robust LMs and for each pair of granularities $(A,B)$, we then calculate the proportion of facts known at granularity $A$, given that they are known at granularity $B$.

### v6

@@ Abstract
This paper explores the robustness of language models (LMs) to variations in the temporal context within factual knowledge. It examines whether LMs can correctly associate a temporal context with a past fact valid over a defined period, by asking them to differentiate correct from incorrect contexts. The LMs’ ability to distinguish is analyzed along two dimensions: the distance of the incorrect context from the validity period and the granularity of the context. To this end, a dataset called TimeStress is introduced, enabling the evaluation of 18 diverse LMs. Results reveal that the best LM achieves a perfect distinction for only 11% of the studied facts, with errors, certainly rare, but critical that humans would not make. This work highlights the limitations of current LMs in temporal representation.
@@ 1 Introduction
When a Language Model (LM) completes the textual prompt "The capital of France is" with "Paris", it demonstrates that it has stored this fact somewhere in its parameters. However, as shown by numerous studies Elazar et al. (2021); Dong et al. (2023); Hagen et al. (2024); Kassner and Schütze (2020), this type of factual knowledge is not necessarily robust to certain variations in the prompt (use of paraphrases, aliases, typographical errors, negations, etc.). Among these variability factors, the temporal dimension of factual knowledge has been less studied. Thus, in this paper, we study the robustness of LMs’ factual knowledge in the face of simple variations in the temporal context.
While the state of the art has demonstrated certain biases in LMs related to the temporal distribution of their training data or their weaknesses in reasoning with temporal concepts, our work aims to quantify how well LMs can correctly associate a temporal context (e.g., a year or a date, such as "In 2018, …", "On November 5, 2022, …") with a past fact, that is, a fact with a certain period of validity. More specifically, the research questions addressed are:
@@
    Do LMs distinguish between correct and incorrect temporal contexts for facts?
@@
Figure 1: The robustness of the LM on a fact is evaluated by asking it to differentiate a set of correct and incorrect statements. The temporal context is varied along two dimensions: its position on the timeline (rows 1 and 2) and its granularity (rows 1 and 3). The trophy means that the sentence was preferred by the LM.
@@
  • The release of a dataset, TimeStress, consisting of popular factual knowledge (according to a popularity index), temporally annotated, and their corresponding high-quality verbalizations. This dataset allows for the replication of our experiments but also opens avenues for other studies on temporality.
  • Highlighting the low robustness of current LMs regarding their factual knowledge when it comes to positioning them in time, as well as errors—certainly rare but critical—that a human would not make. These results reveal the shortcomings of LMs in terms of internal representation of temporality, including for large models (18 models tested across various sizes and families).
In the following sections, we first discuss related work (Section 2). Then, we elaborate on the paper’s issues and present the TimeStress dataset (Section 3). Finally, we describe our experiments and analyze their results (Section 4). The source code and data to reproduce our results will be published soon. The source code enabling the reproduction of our experiments is published on GitHub111github.com/Orange-OpenSource/TimeStress (MIT License) and TimeStress is distributed in Hugging Face222huggingface.co/datasets/Orange/TimeStress (CC BY-SA 4.0 License).
@@ 2 Related work
Several studies have examined the temporal reasoning capabilities of LMs Zhang and Choi (2021); Chu et al. (2024); Wei et al. (2023); Fatemi et al. (2025); Dhingra et al. (2022); Xiong et al. (2024); Su et al. (2024). Notably, the works of Chen et al. (2021) and Tan et al. (2023) each proposed a dataset in which LMs are invited to answer questions involving the understanding of the temporality of facts. While these studies share similarities with ours in terms of data (temporally annotated facts), their objectives and methodologies differ. These studies test the mastery of certain temporal logic operators (date calculations, comparisons, etc.) and evaluate the average performance of LMs based on a one-test-per-fact principle. In contrast, we focus not on reasoning ability but on the robustness of knowledge, that is, the ability of an LM to recall the same fact across various temporal contexts.
@@ 3.1 Problem Statement
Considering a temporal fact $f=(s,r,o,a,b)$, a temporal context $\tau$ is said to be correct for $f$ if $\tau$ is fully included in $[a,b]$ (i.e., $\tau\subseteq[a,b]$), incorrect if it is not included at all ($\tau\cap[a,b]=\varnothing$), or transitional otherwise ($\tau\cap[a,b]\neq\varnothing$ and $\tau\not\subseteq[a,b]$). For example, given the validity period $[2017,2019]$, $2016$ is incorrect, $2017$ is transitional, and $2018$ is correct.
@@
We introduce two metrics. Given a fact $f$ and a model $M$, we express the results using a win rate $\mathcal{W}(M,f)\in[0,1]$ of $M$ for $f$, which is the ratio of the number of times the model preferred a correct context over an incorrect context for the single fact $f$ to the number of tests performed. Additionally, a robustness metric, denoted $\mathcal{R}(M,f)$, verifies that correct contexts consistently outperform incorrect ones, defined as: $\mathcal{R}(M,f)=\mathbbm{1}[\mathcal{W}(M,f)=1]$ where $\mathbbm{1}[]$ is the indicator function. It is important to note that transitional contexts are not used in any way for the calculation of these metrics, as their validity is ambiguous. Given a set of facts, the average win rates and average robustness are denoted $\mathcal{V}(M)$ and $\mathcal{R}(M)$ respectively.
For segmentation purposes in the analyses, these global metrics can be restricted to tests conducted with temporal contexts of a specific granularity (Y, YM, or YMD).
@@ 3.2 The TimeStress Dataset
We present the TimeStress dataset, which enables our study. This dataset contains over 521,000 statements (in the form of questions) generated from 2,003 temporal facts, covering 1,883 unique entities (1,385 unique subjects and 1,113 unique objects) and 86 relations. A brief sample is provided in Table 3.2.3.
On average, each fact is associated with $11$ correct temporal contexts and $74$ incorrect ones, distributed across the three granularities Y, YM, and YMD. There are enough correct and incorrect contexts to make it nearly impossible for a random model to be robust on any fact by chance.
@@ 3.2.1 Quintuplet Collection
The quintuplet collection process begins with a preprocessed version of Wikidata provided in Ammar Khodja et al. (2025). This source also provides a measure of each entity’s popularity, defined as the median number of human visits per month to the Wikipedia article associated with the entity in 2020. This measure is used to define the popularity index of a quintuplet, calculated as the geometric mean of the popularity of its object and subject. Although the popularity of the subject and object does not imply the popularity of the fact, this index remains an interesting tool for finding facts "known" by LMs, as it is shown empirically in the experiments.
We collect and filter Wikidata facts following this procedure: (1) All quintuplets with a validity period (i.e., a start or end date mentioned) and whose objects are not literals, such as quantities and dates, are collected. (2) Quintuplets valid within two distinct periods are removed to simplify result analysis, as this allows all dates outside the validity period to be considered incorrect. (3) Quintuplets without a delimited validity period (i.e., a start AND end date mentioned) are removed. (4) Only quintuplets that were valid prior to 2021 are retained, as this ensures that all these quintuplets are past facts for all studied LMs. (5) Only the quintuplets that are valid for longer than three years are retained to ensure a minimal number of correct temporal contexts of Y granularity. (6) We keep only the most popular quintuplets using the popularity index. This results in a set of 2,098 quintuplets with a varied set of 86 relations.
@@ 3.2.2 Quintuplet Verbalization
The process of generating statements from quintuplets is carried out using GPT-4o. First, a prompt instructs GPT-4o to generate four linguistically diverse questions from a given tuple (subject, relation, object, year), with the following guidelines: the question must be in the past tense, begin with “In [YEAR],”, be stated in a simple and concise manner without any detail that could give clues about the answer. It should be directly followed by the answer, which is the object. The quality of the generated questions was analyzed to identify and eliminate incorrect entries. Initially, out of the 2,098 facts intended for verbalization, 53 failed, and 64 questions mistakenly used the subject as the answer instead of the object. These erroneous cases were removed from the dataset, resulting in a total of 2,003 facts and $2003\times 4=8012$ questions. A random sample of 50 questions was manually evaluated to ensure the overall quality of the generated questions. The evaluation revealed that only 1 out of 50 questions was incorrect, while the remaining questions were perfectly constructed (Wilson confidence interval at 95% = [0.85, 0.99]), which demonstrates the high quality of the questions in our dataset. Finally, the temporal context was removed and each fact is randomly assigned one of its four associated questions.
@@ 3.2.3 Context Sampling
For each fact, based on its validity interval $[a,b]$, centered on $m=\frac{a+b}{2}$ and of duration $d=b-a$ [^3] The median of dates (in day precision) is used to perform arithmetic operations between dates., temporal contexts at the Y granularity are uniformly sampled over the wider interval $[m-5d,m+5d]$ with a step of $0.05\times d$. From these Y-granularity contexts, YM-granularity contexts are generated by randomly selecting a month. Similarly, YMD-granularity contexts are determined by choosing a random day from each YM-granularity context [^4] This sampling does not produce erroneous dates such as February 29 for non-leap years, or April 31.. This process creates a hierarchy among contexts derived from the same year for a given fact. Note that when a date $d_{2}$ is chosen from a higher-granularity date $d_{1}$, it is necessarily correct (or incorrect) if $d_{1}$ is. However, $d_{1}$ may be transitional while $d_{2}$ is correct or incorrect. In such cases, $d_{2}$ is excluded from the set of correct or incorrect dates. This guarantees that the number of correct and incorrect contexts does not vary by granularity, avoiding bias when comparing model robustness across granularities. The corresponding years of the produced contexts are mainly located in the contemporary period between 1800 and 2020 (Appendix E), because the popularity index used to select the facts in TimeStress draws more often recent facts. To produce the statements of each fact that will be used to compute the metrics, its corresponding statement is prefixed with the previously sampled temporal contexts associated with the fact.
@@ 4 Experimentation
This section details our experiments on the TimeStress dataset. As a reminder, our objectives are, in order, to measure the ability of models to distinguish correct and incorrect temporal contexts, analyze their robustness, and search for anomalies in this task when incorrect contexts are closer to or farther from the validity interval, and as the granularity of contexts becomes finer.
Numerous models from different families and sizes were tested: Mistral-Nemo-Base-2407, Mistral-7B-v0.3 Jiang et al. (2023); OpenEML-{450M, 3B} Mehta et al. (2024); gemma-2-{2b, 9b, 27b} Team et al. (2024); Llama-3.1-{8B, 70B} Grattafiori et al. (2024). For each, both pretrained and instruction-tuned versions were considered, resulting in a total of 18 studied LMs. All models were sourced from huggingface.co.
In the first series of experiments, the statements were passed to the models as raw text rather than as instructions to enable the comparison between pretrained and instruction-tuned models. The use of an "instruction/message" format is explored in a second phase.
@@ 4.1 Overall Mastery of Temporal Contexts
Figure 2(a) shows the average win rate for the facts in TimeStress for the top 5 LMs and for each temporal granularity Y, YM, and YMD, as well as for their union. Results for other models are reported in Appendix E.
Overall, the results show that these top 5 LMs generally distinguish correct statements from incorrect ones well, with win rates ranging from 78% to 87%. Among our other findings, we observed that even smaller models (<500M parameters) perform better than chance, and the win rate logically improves with model size (Appendix E), with the best model being the largest, Llama-3.1-70B-Instruct.
Figure 3 provides a more detailed analysis by reporting the average $\log\Pr(o|f,\tau)$ as a function of the value $\alpha$, which quantifies the relative distance of $\tau$ from the validity period of $f$ (see Section 3.1). The average is calculated across all facts, for contexts at the year granularity, and across all 18 studied LMs. We observe that the highest probabilities correspond to contexts within the validity interval ($\alpha\in[-0.5,0.5]$), while outside this interval, probabilities gradually decrease as $|\alpha|$ increases. Finally, we note that the probability assigned to transitional contexts (years that are neither fully correct nor fully incorrect) is significantly higher (based on the confidence intervals (CIs)) than that for incorrect contexts. We explain this phenomenon with the following hypothesis: in the training data of LMs, transitional years are more often associated with the considered fact than other years within the validity period, as they correspond to key events such as the beginning and end of the fact (e.g., the start or end year of a presidential term).
@@
Figure 2: Average metrics on the TimeStress dataset for the 5 most robust models (95% CIs were determined using bootstrapping).
@@ 4.2 Robustness and Anomalies
Figure 2(b) shows the average robustness of the top 5 models across all facts in TimeStress. As a reminder, this metric is stricter and does not tolerate any error during matches for a given fact. Results for other models are reported in Appendix E.
The scores are generally low, indicating that win rates per fact rarely reach 100%. Interestingly, the most robust model is not the one with the highest win rate. The most robust model, gemma-2-27b-it, achieves an $\mathcal{R}$ value of only about 17% for the coarsest granularity Y. This score drops to 11% when all granularities are considered. Most other models do not exceed a global robustness score of 3%. Among our other results, we also observed that instruction-tuned models mostly outperform their pre-trained counterparts. A notable case is the Llama-3.1-70B-Instruct model; although it was fine-tuned on instructions, it is 3.6×3.6$\times$3.6 × more robust than its pre-trained counterpart, Llama-3.1-70B. This suggests that the training data and possibly the training procedure play an important role in temporal robustness. Finally, early signs of failure in knowledge transfer between granularities are evident due to the substantial gap between individual robustness scores for granularities and the global score. This issue is explored in detail later in this section.
Figure 3: Evolution of $\log\Pr(o|f,\tau)$ with respect to the relative distance $\alpha$, averaged across all facts in TimeStress and all LMs, for granularity Y (Bootstrap 95% CIs). The number of points used to compute each bar is indicated above it.
$|\alpha|$ Raw Text Instruction
$\geq 1$ 0.19 (±0.01) 0.25 (±0.01)
$\geq 2$ 0.09 (±0.01) 0.13 (±0.01)
$\geq 3$ 0.06 (±0.01) 0.08 (±0.01)
$\geq 4$ 0.04 (±0.01) 0.05 (±0.01)
Figure 4: Proportion of incorrect dates favored over correct dates beyond a relative distance $|\alpha|$, when the win rate exceeds 95% (Wilson’s 95% CIs).
@@
Table 4 investigates the impact of the relative positions of incorrect contexts of granularity Y, focusing on cases where incorrect contexts cause an LM to fail in a match for facts that seem "known" to the LM, as indicated by a very high win rate ($\mathcal{W}\geq$ 95%). For now, only the "Raw Text" column is of interest. The table reveals that these incorrect contexts are not entirely concentrated around the validity period, as might reasonably be expected. Instead, a significant proportion is located far from it. Specifically, LMs fail to achieve robustness due to contexts with a distance of $|\alpha|\geq 1$ in 19% of cases. This proportion decreases to 6% for $|\alpha|\geq 3$, which remains significant given the proximity of the win rate to 100% for the facts observed here. We conducted the same analysis using win rate thresholds higher than 95% (see Appendix B). As the threshold approaches 100%, vulnerability to "easy" incorrect dates gradually decreases but never completely disappears. Even when the win rate threshold is 99%, errors remain when $|\alpha|\geq 4$. We conclude that this vulnerability is inherent to current LMs. While the probabilistic nature of these models may provide a tangible explanation, this behavior is clearly undesirable, as these are typically errors that a human would not make when aware of a fact’s validity period.
@@
So far in our experiments, all models have been fed statements in Raw text rather than instructions. Since the performance of instruction-tuned LMs might have been underestimated, win rates and robustness scores were recalculated using an "instruction/message" format [^5] This involves constructing messages and injecting them into the chat template of each LM, as in the following example: {user: "In 2011, who was the president of the USA?", assistant: "Barack Obama"}.. Figure 5 compares robustness scores calculated for the two formats. On average, robustness decreases with the use of the "instruction" format (notably for gemma-2 models), and global robustness scores remain low. However, no clear conclusions emerge regarding the positive or negative impact of this format, as the effect varies significantly across models. Next, the "Instruction" column of Table 4 complements our previous analysis on the impact of the relative position of incorrect contexts for high win-rate facts. This time, the "instruction" format degrades performance with more critical errors (i.e., far from the validity period). Based on the confidence intervals, these differences are statistically significant for all values of $|\alpha|$ studied. Examples of these critical errors are shown in Appendix E.
@@
Figure 5: Average $\mathcal{R}$ across all granularities for facts in TimeStress based on the format of statements submitted to the models: raw text (blue) or instruction (orange). 95% CIs were determined using bootstrapping.
@@
Figure 6: Average success rate of knowledge transfer between granularity pairs for the 5 most robust LMs with queries in raw text (left) or instructions (right). Wilson confidence intervals at 95% are shown.
LMs fail to perfectly propagate their knowledge across granularities.
@@
We examine the ability of LMs to propagate knowledge of a fact across different temporal granularities. TimeStress allows comparisons between two granularities because the three studied granularities have the same number of correct and incorrect contexts for all temporal facts. The only difference between two granularities is the addition of a random month and/or day, which does not affect validity when transitioning from a lower granularity to a higher granularity (e.g., from Y to YM). For example, if a fact is incorrect for an entire year, it remains incorrect for any month or date within that year.
@@
We consider a fact $f$ to be "known" for a granularity by a model $M$ if $\mathcal{R}(M,f)=1$. This definition can apply to a given granularity. For example, a fact is "known" at the Y granularity if all matches with temporal contexts at the year granularity were won. For each of the 5 most robust LMs and for each pair of granularities $(A,B)$, we then calculate the proportion of facts that are "known" at granularity $A$, given that they are "known" at granularity $B$.